OpenStudy (fantageplayer):

The lengths of two sides of a triangle are shown below: Side 1: 3x2 − 2x − 1 Side 2: 9x + 2x2 − 3 The perimeter of the triangle is 5x3 + 4x2 − x − 3. Part A: What is the total length of the two sides, 1 and 2, of the triangle? (4 points) Part B: What is the length of the third side of the triangle? (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer. (2 points)

OpenStudy (welshfella):

Part A Just add like terms :- 3x^2 - 2x - 1 + 9x + 2x^2 - 3

OpenStudy (fantageplayer):

this is what i have so far: Part A: ](3x2 − 2x − 1)+(9x + 2x2 − 3) 3x2−2x−1+9x+2x2−3 =3x2+−2x+−1+9x+2x2+−3 Combine Like Terms: =3x2+−2x+−1+9x+2x2+−3 =(3x2+2x2)+(−2x+9x)+(−1+−3) =5x2+7x+−4 Answer: =5x2+7x−4 Part B:] 5x3+4x2−x−3−(5x2+7x−4) Distribute the Negative Sign: =5x3+4x2−x−3+−1(5x2+7x−4) =5x3+4x2+−x+−3+−1(5x2)+−1(7x)+(−1)(−4) =5x3+4x2+−x+−3+−5x2+−7x+4 Combine Like Terms: =5x3+4x2+−x+−3+−5x2+−7x+4 =(5x3)+(4x2+−5x2)+(−x+−7x)+(−3+4) =5x3+−x2+−8x+1 Answer: =5x3−x2−8x+1

OpenStudy (welshfella):

Par5 B subtract the result from the above addition from the expression for the perimeter.

OpenStudy (fantageplayer):

yah i just dont knoow what part c is supposed to mean

OpenStudy (welshfella):

yes those 2 parts are correct

OpenStudy (fantageplayer):

thanks but do you know what part c means?

OpenStudy (welshfella):

polynomials being closed under addition and subtraction means that if you add or subtract polynomials then the result is another polynomial.

OpenStudy (welshfella):

so in your case do the results support this?

OpenStudy (fantageplayer):

yeah

OpenStudy (welshfella):

yea they do Thats all there is to Part C.

OpenStudy (fantageplayer):

thanks!

OpenStudy (fantageplayer):

thanks!

OpenStudy (fantageplayer):

Part C: Yes, when i added side 2 and side 1 i got 5x2+7x−4 which is another plynomial

OpenStudy (fantageplayer):

does that sound good @welshfella